Find class boundaries, midpoint, and width for the class. 121-133
Answer #1

Class width = 134-121 =13

Midpoint = lower class limit+upper class limit/2 = 121+133/2=127

Midpoint is the centre/average of class limits

Class boundries are use to make distribution perfectly continious

Class bounderies = 120.5-133.5